OpenStudy (anonymous):

Some help needed. This is about writing function rules for situations so here are my questions... the height f(h) of an object in feet when you know the height h in inches. the volume V(n) of a cube when you know the length n of a side.

OpenStudy (phi):

the height f(h) of an object in feet when you know the height h in inches. what rule would you use to find the number of feet in 12 inches?

OpenStudy (anonymous):

idk, it just says in the directions to write a function rule.

OpenStudy (phi):

If you measure 12 inches how many feet is that?

OpenStudy (anonymous):

1

OpenStudy (phi):

if you measure 24 inches, how many feet?

OpenStudy (anonymous):

2

OpenStudy (phi):

how are you getting your answers of 1 and 2? did you divide by 12 by any chance?

OpenStudy (anonymous):

i kinda just knew them,but i see that your trying to point out that this should probabaly be the answer f(h)=h divided by 12,right? ok. So how about the second one?

OpenStudy (phi):

yes f(h)= h/12 is how I would write it.

OpenStudy (phi):

the volume V(n) of a cube when you know the length n of a side. you need to know how to find the volume of a cube. Google works, but *everybody* knows that volume = n^3 (n*n*n is another way to write it)

OpenStudy (anonymous):

so... v(n)=n^3?

OpenStudy (phi):

yes.

OpenStudy (anonymous):

Omg,thanks sooo much. You make ths easy to understand and you dont overcomplicate things!